OpenStudy (anonymous):

Solve the following equation. 96=-6(-4-3y) A.y=2 B.y=4 C.y=-2 D. y=-4 WILL GIVE MEDAL FOR CORRECT ANSWER!!!

OpenStudy (anonymous):

B, y = 4.

OpenStudy (anonymous):

-4 - 3 x 4 = -16 -16 x -6 = 96